OpenStudy (anonymous):

Hi all , I am trying to solve for x sqaure root of 9x^2 - 16 = 3x - 4 can anyone show me the wroking out for this? Much appreciated A

OpenStudy (anonymous):

\[\sqrt{9x^2 - 16} = 3x - 4\] Is this what you mean?

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

Square both sides and rearrange, your x^2 terms will cancel.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

9x^2 - 16 = (3x - 4)^2 9x^2 - 4^2 = 9x^2 + 4^2 - 2* 3* 4*x 2*4^2 = 2 * 3 * 4 * x 4 = 3 * x x = 4/3